OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length sqrt180 of feet and a width of sqrt125

OpenStudy (anonymous):

perimeter= (L+w)2 = (sqrt180+sqrt125)2 = 49.19

OpenStudy (anonymous):

could you go into more detail plz, this is the only question i cant figure out

OpenStudy (anonymous):

the formula is perimeter = (length + width) times 2

OpenStudy (anonymous):

this is multiplying and dividing radicals

OpenStudy (anonymous):

p= (l+w)2 =(6sqrt5 +5sqrt5)2 =11sqrt5 times 2 =22sqrt5

OpenStudy (anonymous):

\[\sqrt{180}=\sqrt{9*4*5}=6*\sqrt{5}\]

OpenStudy (anonymous):

Do same thing to sqrt of 125

OpenStudy (anonymous):

sqrt(25*5)